how many ways can thirteen objects be selected four at a time
Answer #1

The no of ways thirteen objects be selected four at a time

= _{4}^{13}\textrm{C}

= \frac{13!}{(13-4)!*4!}

= \frac{13!}{9!*4!}

\frac{13*12*11*10*9!}{9!*4!}

\frac{13*12*11*10}{24}

= 715

So, answer is 715